OpenStudy (anonymous):

What sum of money invested at 7.5% per annum will yield same simple interest in 4 years as rs5000 yields in 5 years at 9% per annum?

OpenStudy (wolf1728):

We'll have to answer the compound interest problem first. total = princ*(1+rate)^years total = 5,000 * 1.09 ^5 total = 5,000 * 1.5386239549 total = 7,693.12 For the simple interest problem, we need to know the money needed to be invested for 5 years to yield $7,693.12. princ + .09 * 5 * princ = 7,693.12 princ + .45*princ = 7,693.12 princ. = 7,693.12 / 1.45 = 5305.60 5305.60*.09 = 477.50 interest per year * 5 years = 2,387.50 5,305.60 plus 2,387.50 equals 7,693.10 Total